# FMDLlama: Financial Misinformation Detection based on Large Language Models

### v1

@@ 5 Conclusion
In this paper, we propose FMDLlama, the first LLM for financial misinformation detection (FMD). We also construct a multi-task FMD instruction dataset (FMDID) and a FMD evaluation benchmark (FMD-B). We conduct a comprehensive analysis of the performance of FMDLlama, as well as a variety of LLMs on the FMD-B benchmark. The results indicate that FMDLlama performs exceptionally well in FMD tasks, achieving SOTA compared to the other open-sourced LLMs as well as ChatGPT.
@@ 6 Limitations
(1) Due to restricted computational resources, we only carried out instruction-tuning/evaluation of financial misinformation detection tasks using 7b/13b LLMs. As such, we have not considered the impact of using larger models on the FMD tasks.

### v2

@@ 5. Conclusion
In this paper, we propose FMDLlama, the first LLM for financial misinformation detection (FMD). We also construct a multi-task FMD instruction dataset (FMDID) and a FMD evaluation benchmark (FMD-B). We conduct a comprehensive analysis of the performance of FMDLlama, as well as a variety of LLMs on the FMD-B benchmark. The results indicate that FMDLlama performs exceptionally well in FMD tasks, achieving SOTA compared to the other open-sourced LLMs as well as OpenAI’s products.
@@ 6. Limitations
(1) Due to restricted computational resources, we only carried out instruction-tuning/evaluation of financial misinformation detection tasks using 2b/7b/8b/13b LLMs. As such, we have not considered the impact of using larger models on the FMD tasks.
@@
Acknowledgements.
The FMDLlama illustration in Figure 1 was generated using PIXLR [^7] https://pixlr.com/image-generator/. This work is supported by the computational shared facility at the University of Manchester and the scholar award from the Department of Computer Science at the University of Manchester. This work is supported by the project JPNP20006 from New Energy and Industrial Technology Development Organization (NEDO) and the Manchester-Melbourne-Toronto Research Funding. This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGeneration EU Program.